### v2

# Key-Point-Driven Mathematical Reasoning Distillation of Large Language Model

Abstract
Large Language Models (LLMs) have demonstrated exceptional proficiency in mathematical reasoning tasks due to their extensive parameter counts and training on vast datasets. Despite these capabilities, deploying LLMs is hindered by their computational demands. Distilling LLM mathematical reasoning into Smaller Language Models (SLMs) has emerged as a solution to this challenge, although these smaller models often suffer from errors in calculation and semantic understanding. Prior work has proposed Program-of-Thought Distillation (PoTD) to avoid calculation error. To further address semantic understanding errors, we propose Key-Point-Driven Mathematical Reasoning Distillation (KPDD). KPDD enhances the reasoning performance of SLMs by breaking down the problem-solving process into three stages: Core Question Extraction, Problem-Solving Information Extraction, and Step-by-Step Solution. This method is further divided into KPDD-CoT, which generates Chain-of-Thought rationales, and KPDD-PoT, which creates Program-of-Thought rationales. The experiment results show that KPDD-CoT significantly improves reasoning abilities, while KPDD-PoT achieves state-of-the-art performance in mathematical reasoning tasks. Our approach effectively mitigates misunderstanding errors, advancing the deployment of efficient and capable SLMs.
Key-Point-Driven Mathematical Reasoning Distillation of Large Language Model
Xunyu Zhu1,2, Jian Li1,2thanks: Corresponding author, Can Ma1,2, Weiping Wang1,2 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences {zhuxunyu, lijian9026, macan, wangweiping}@iie.ac.cn
1 Introduction
Large language models (LLMs) have achieved impressive performance in mathematical reasoning tasks. Recent work (Wang et al., 2023) further proposes Chain-of-Thought (CoT) to enhance the mathematical reasoning abilities of LLMs. However, the massive scale of LLMs presents significant challenges for deployment.
Figure 1: Error analysis of 50 GSM8K problems with incorrect answers returned by CoTD using FlanT5-Base. The experimental results indicate that multiple errors may exist in the reasoning process of CoTD, with understanding errors and calculation errors being the major factors affecting CoTD’s reasoning performance.
A feasible solution to address this problem is to use black-box distillation to transfer mathematical reasoning abilities from LLMs to SLMs. Chain-of-Thought Distillation (CoTD) Ho et al. (2023) is a representative mathematical reasoning distillation method. CoTD prompts LLMs to generate reasoning processes for each question, constructing a mathematical reasoning dataset. This dataset is then used to fine-tune SLMs, enhancing their mathematical reasoning abilities. However, there remains a significant performance gap between SLMs and LLMs. Prior work (Wei et al., 2022) identifies three main error types in CoT reasoning: Calculation errors, Missing Step errors, and Semantic misunderstanding errors. To explore the reasons for the performance gap between SLMs and LLMs, we conducted the same error analysis on CoTD. Our preliminary experiments (shown in Figure 1) reveal numerous error combinations in CoTD, with calculation and semantic misunderstanding errors being the most prevalent. Prior work Zhu et al. (2023) proposed Program-of-Thought Distillation (PoTD) to avoid calculation errors by formulating the reasoning process as a Python program executed by an external interpreter. This approach allows the SLM to focus on generating the program, avoiding calculation errors and improving reasoning performance. Given these circumstances, our paper focuses on addressing semantic misunderstanding errors in CoTD to further enhance the reasoning performance of SLMs.
In our paper, we propose a novel mathematical reasoning distillation method called Key-Point-Driven Mathematical Reasoning Distillation (KPDD) to enhance the mathematical reasoning performance of SLMs. KPDD breaks the reasoning process into three parts: (1) Core Question Extraction: Identifies the core question from the original problem. (2) Problem-Solving Information Extraction: Extracts relevant data and information needed to solve the problem. (3) Step-by-Step Solution: Uses the extracted key points to solve the problem in a step-by-step manner. The third part is further divided into two formats, KPDD-CoT and KPDD-PoT: (1) KPDD-CoT: Generates rationales in the form of Chain-of-Thought (CoT). This method focuses on reducing misunderstanding errors and explicitly illustrates the reasoning process, aiding in error analysis. (2) KPDD-PoT: Generates rationales in the form of Program-of-Thought (PoT). This approach not only reduces misunderstanding errors but also avoids calculation errors, further enhancing the SLM’s mathematical reasoning performance.
We use KPDD to fine-tune FlanT5 models, and evaluate these models on several mathematical reasoning dataset, including GSM8K, ASDiv, SVAMP, and MultiArith. Our experiment results show that KPDD-CoT significantly enhances SLMs’ reasoning abilities, while KPDD-PoT enables SLMs to achieve state-of-the-art (SOTA) mathematical reasoning performance. Furthermore, our error analysis on KPDD confirms that KPDD effectively mitigates misunderstanding errors, thereby improving the mathematical reasoning performance of SLMs.
Our contributions are summarized as follows:
  1. 1.
    Our study reveals that misunderstanding errors and calculation errors are the major factors limiting CoTD’s reasoning.
  2. 2.
    We propose Key-Point-Driven Mathematical Reasoning Distillation (KPDD) to alleviate misunderstanding errors and effectively improve the reasoning performance of SLMs.
  3. 3.
    Extensive experiments show that KPDD outperforms other methods across various benchmarks and achieves new state-of-the-art results on these mathematical reasoning datasets.
2 Related Work
2.1 Chain-of-Thought Reasoning
Chain-of-Thought refers to prompt LLMs to solve the question step by step. Prior work Wei et al. (2022) find that Chain-of-Thought can effectively improve the reasoning performance of LLMs. Based on the findings, Kojima et al. Kojima et al. (2022) further introduce zero-shot CoT, which significantly improves the model’s reasoning performance by simply adding the prompt "Let’s think step by step" before answering. To avoid calculation error in CoT, Chen et al. Chen et al. (2023) formulate the reasoning process into program. Furthermore, Wang et al. Wang et al. (2023) introduce a self-consistency decoding strategy, which generates diverse reasoning paths and then selects the most consistent answer by considering these paths comprehensively. Least-to-most prompting Zhou et al. (2023) breaks down complex problems into a series of simpler subproblems and solves these subproblems sequentially. Zhong et al. Zhong et al. (2024) encourage LLMs to deeply understand problems and leverage key information for better reasoning. Inspired by these methods, our work introduces Key-Point-Driven Mathematical Reasoning Distillation (KPDD) to further enhance SLMs’ mathematical reasoning.
2.2 Black-box Distillation
Currently, close-source LLMs usually have stronger reasoning performance than open-source LLMs. However, in general, we can only obtain the output of close-source LLMs. Based on this situation, black-box distillation is proposed to distill abilities from LLMs to SLMs. Specifically, black-box distillation first prompts LLMs to generate a distillation dataset, and then this dataset is used to fine-tune SLMs to improve their reasoning performance. For example, Ho et al. Ho et al. (2023) prompt LLMs to generate a CoT reasoning dataset, which was then used to fine-tune an SLM, thereby enhancing its reasoning ability. Shridhar et al. Shridhar et al. (2023) utilize a LLM to train a problem decomposer and a subproblem solver separately. They used the decomposer to break down the original problem into multiple subproblems, and then employed the subproblem solver to solve each subproblem individually. By integrating the solutions to these subproblems, they derived the final answer. Fu et al. Fu et al. (2023) analyze CoT distillation and find that there is a trade-off between the specific abilities and the general abilities of SLMs. When CoT distillation is used to enhance the specific abilities of SLMs, their general abilities decrease correspondingly. Zhu et al. Zhu et al. (2023) construct a distillation dataset, where the reasoning format changed from CoT reasoning to PoT reasoning. Through this method, the SLM can delegate the calculation process to an additional Python interpreter to avoid calculation errors, thereby enhancing the SLM’s mathematical reasoning performance. Zhu et al. Zhu et al. (2024) further formalize the reasoning process into equations and find that a diverse range of reasoning formats can effectively enhance the mathematical reasoning performance of SLMs. Our work introduces a novel distillation approach where two SLMs independently extract the core question and key problem-solving information from an original question. These key points are then utilized to guide another SLM in solving the original question effectively.
3 Method
Figure 2: Detailed data generation for KPDD-CoT. Initially, we use few-shot prompting to guide LLMs in producing reasoning processes. We then extract the correct answers from these processes, discarding any reasoning process that does not align with the correct answers. In this way, we have constructed a high-quality KPDD-CoT reasoning distillation dataset.
In this work, we introduce a novel distillation method for mathematical reasoning tasks called Key-Point-Driven Distillation (KPDD), structured into three stages: (1) Stage 1: KPDD distills the first SLM to extract the core question from the original question. (2) Stage 2: KPDD distills the second SLM to extract problem-solving information from the original question. (3) Stage 3: KPDD distills the third SLM to solve the original problem using the core question and problem-solving information. In Stage 3, we prompt the LLM to construct two types of reasoning datasets: (1) CoT Rationales: These are more comprehensible to both humans and LLMs, showcasing a detailed reasoning process. (2) PoT Rationales: These rationales delegate computational tasks to an external Python interpreter, thereby avoiding calculation errors.
3.1 Data Generation from LLMs
When given a mathematical dataset, our primary task is to utilize an LLM to generate a reasoning process for each mathematical problem in the dataset. By doing so, we augment the mathematical dataset to construct the distillation dataset. Furthermore, in stage 3, our KPDD method employs two distillation approaches: one distills the SLM to generate CoT rationales for problem-solving, and the other distills the SLM to generate PoT rationales for problem-solving. In other words, our KPDD method can be divided into two approaches: KPDD-CoT and KPDD-PoT.
3.1.1 Data Generation for KPDD-CoT
Our method uses few-shot prompting to prompt the LLM to synthesize the reasoning dataset. Figure 2 reveals the distillation dataset generation process. Specifically, we first randomly sample $k$ data pairs $(x,y)$ from a mathematical dataset $\mathcal{D}$, where $x$ represents the question and $y$ represents the answer. Then, for these sampled data, we manually construct the reasoning process $c$. Each reasoning process $c$ includes a core question, problem-solving information, and rationales in CoT format. These elements are separated by HTML tags: "<core>{core question}</core><info>{problem-solving information}</info><cot>{rationales in CoT format}</cot>". Finally, we obtain a demonstration dataset $\mathcal{D}_{c}$. At the same time, we further introduce an instruction "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, solve the question step by step, and show the answer.". By leveraging the demonstration set and the instruction, we prompt the LLM to generate the reasoning process for the mathematical questions. The KPDD-CoT dataset generation is formalized as:
$c_{i}=f_{\mathcal{M}}(x_{i},\mathcal{D}_{c}),$ (1)
where $\mathcal{M}$ denotes the LLM, $f$ represents the decoding function, and $i$ is the index in $\mathcal{D}$.
Data Filtering—To further improve the quality of the dataset, we attempt to filter out incorrect reasoning processes from the KPDD-CoT dataset. Specifically, for each data point in the KPDD-CoT dataset, we extract the answer from its reasoning process. If this answer matches the gold answer, we retain the data point; otherwise, we discard it as the reasoning process is deemed incorrect. By ensuring that the KPDD-CoT dataset contains only correct reasoning processes, we enhance the LLM’s mathematical reasoning performance.
Figure 3: Detailed data generation for KPDD-PoT. Similar to KPDD-CoT, we prompt LLMs to generate the reasoning process in the KPDD-PoT format. We extract the rationale in the PoT format from this process and run it through a Python interpreter. If there are errors or incorrect answers, we discard the reasoning process. Finally, we constructed a high-quality KPDD-PoT reasoning dataset.
3.1.2 Data Generation for KPDD-PoT
Similar to KPDD-CoT, we first sample $k$ examples from a mathematical dataset $\mathcal{D}$, then use these samples to manually create a demonstration set $\mathcal{D}_{p}$. Each reasoning process in the demonstration set $\mathcal{D}_{p}$ includes a core question, problem-solving information, and rationales in PoT format. These elements are also separated by HTML tags: "<core>{core question}</core><info>{problem-solving information}</info><pot>{rationales in PoT format}</pot>". Then, we further introduce the instruction a instruction "Firstly, let’s extract the most comprehensive and detailed key question. Then, let’s identify and list the most useful information related to the question. Finally, let’s understand the key question and the problem-solving information, and generate the python code (return ans) to solve the question.". By utilizing the demonstration set $\mathcal{D}_{p}$ and the instruction, we prompt LLMs to generate reasoning processes in for questions. Figure 3 shows the data synthesis process for KPDD-PoT, and the KPDD-PoT dataset generation process can be formalized as:
$p_{i}=f_{\mathcal{M}}(x_{i},\mathcal{D}_{p}).$ (2)
Data Filtering—Similar to data filtering for KPDD-CoT, we also apply filtering to the KPDD-PoT dataset. Specifically, for each data point in the KPDD-PoT dataset, we extract the rationale in PoT format from its reasoning process. An external Python interpreter is then used to run the rationale. If the obtained answer do not match the gold answer, it indicate that the reasoning process is incorrect. Consequently, we remove the data point with the flawed reasoning process from the KPDD-PoT dataset, thereby enhancing the dataset’s quality.
3.2 Fine-tuning SLMs
After constructing these reasoning datasets, we use them to fine-tune the SLMs. In the KPDD, we fine-tune three SLMs: the first SLM, called KPDD-CoT/PoT-core, is used to extract the core question from the original problem, the second SLM, called KPDD-CoT/PoT-info, extracts the problem-solving information, and the third SLM, called KPDD-CoT/PoT-solve, uses both the core question and problem-solving information to solve the original question.
3.2.1 Fine-tuning SLMs for KPDD-CoT
Firstly, we construct a core question subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CC}$. Each sample in this subset can be represented as $(x,cc)$, where $x$ represents the original question and $cc$ represents the core question. For each training instance $(x,cc)$ from $\mathcal{D}_{CC}$, we prepend the prompt $p_{cc}$Let’s extract the most comprehensive and detailed core question." to the question $x$. This guides the KPDD-CoT-core in fine-tuning to accurately extract the corresponding core question $cc$. The fine-tuning loss function can be represented as follows:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T}\log P({cc}^{i}_{t}\mid{cc}^{i}_{<t},x^{i},p_{cc}),$ (3)
where $N$ is the number of examples in $\mathcal{D}_{CC}$, $p_{cc}$ is the prompt, and ${cc}_{:T}$ is the sequence of the core question.
Then, we construct a problem-solving subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CI}$. Each sample in this subset can be represented as $(x,ci)$, where $x$ represents the original question and $ci$ represents the problem-solving information. For each data point $(x,ci)$ from $\mathcal{D}_{CI}$, we add the prompt $p_{ci}$Let’s identify and list the most useful information related to the question." to the question $x$. This guides the KPDD-CoT-info in fine-tuning to accurately extract the corresponding problem-solving information $ci$. The fine-tuning loss function can be represented as follows:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T}\log P({ci}^{i}_{t}\mid{ci}^{i}_{<t},x^{i},p_{ci}),$ (4)
where $N$ is the number of examples in $\mathcal{D}_{CI}$, $p_{ci}$ is the prompt, and ${ci}_{:T}$ is the sequence of the problem-solving information.
Figure 4: The inference process of KPDD. When given an original question, KPDD-CoT/PoT-core and KPDD-CoT/PoT-info first extract the core question and the problem-solving information. Then, KPDD-CoT/PoT-solve uses these key points to generate rationales to solve the original question.
Finally, we construct a problem-solving subset from the KPDD-CoT dataset, denoted as $\mathcal{D}_{CS}$. Each sample in this subset can be represented as $(x,cc,ci,cs)$, where $x$ represents the original question, $cc$ represents the core question, $ci$ represents the problem-solving information, and $cs$ represents the rationales in CoT format. For each data $(x,cc,ci,cs)$ from $\mathcal{D}_{CS}$, we integrate the original question $x$, the core question $cc$, the problem-solving information $ci$ and the prompt $p_{cs}$Let’s understand the core question and the problem-solving information, solve the question step by step, and show the answer." to construct a new input. This guides the KPDD-CoT-solve in fine-tuning to generate rationales $cs$ for solving the origin question in CoT format. The fine-tuning loss function can be represented as follows:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{t=1}^{T}\log P({cs}^{i}_{t}\mid{cs}^{i}_{<t},x^{i},cc^{i},ci^{i},p_{cs}),$ (5)
where $N$ is the number of examples in $\mathcal{D}_{CS}$, $p_{cs}$ is the prompt, and ${cs}_{:T}$ is the sequence of the rationale in CoT format.
3.2.2 Fine-tuning SLMs for KPDD-PoT
In KPDD-PoT, aside from replacing the KPDD-CoT dataset with the KPDD-PoT dataset, the fine-tuning method for KPDD-PoT-core remains consistent with that of KPDD-CoT-core, and the fine-tuning method for KPDD-PoT-info remains consistent with that of KPDD-CoT-info. However, the fine-tuning method of KPDD-PoT-solve is different with KPDD-CoT-solve. The main difference between them is the input instruction. Specifically, when fine-tuning KPDD-PoT-solve, the input instruction is: "Let’s understand the core question and the problem-solving information, and generate the python code (return ans) to solve the question." This instruction guides the model to not only understand the core question and the problem-solving information but also to generate Python code that can compute the answer. This approach leverages the model’s ability to perform code generation, which can be particularly effective for solving mathematical problems programmatically.
Moreover, the fine-tuning loss functions for the SLMs in KPDD-PoT are identical to those in KPDD-CoT. This ensures that the optimization process remains consistent across both methods, focusing on minimizing the discrepancies between the model’s output and the expected solutions.
3.3 Inference-time Predictions
Figure 4 illustrates the inference process of KPDD. After fine-tuning, the process for solving a given question involves three main steps:
  1. 1.
    Core Question Extraction: First, we use the KPDD-CoT/PoT-core model to extract the core question from the original problem. This step isolates the essential part of the problem that needs to be addressed.
  2. 2.
    Problem-Solving Information Extraction: Next, the KPDD-CoT/PoT-info model extracts the relevant problem-solving information. This model identifies and lists the necessary context and data required to solve the core question.
  3. 3.
    Solution Generation: Finally, based on the original question, the core question, and the problem-solving information, the KPDD-CoT/PoT-solve model generates rationales in either CoT or PoT format to solve the original question. For KPDD-PoT, this involves generating Python code that can compute the answer.
This structured approach ensures that each model focuses on a specific aspect of the problem-solving process, leading to more accurate and reliable solutions.
4 Experiments
Dataset Size
Train GSM8K 7473
(+) augmented 29892
Test GSM8K 1319
ASDiv 2096
SVAMP 1000
MultiArith 600
Table 1: Statistics of the datasets used in our experiments. Augmented refers that we run 4 times data synthesis on the training set of GSM8K.
4.1 Dataset
In our paper, we generate KPDD distillation datasets based on the GSM8K training set, which comprises diverse grade school math word problems (Cobbe et al., 2021). Then, we evaluate the mathematical reasoning performance of SLM on the GSM8K test set. Furthermore, to assess the transferability of SLMs’ mathematical reasoning capabilities, we evaluate SLM on several additional mathematical datasets. These datasets include ASDiv, which contains diverse math word problems (Miao et al., 2020), SVAMP, which features math word problems with varying structures (Patel et al., 2021), and MultiArith, which consists of arithmetic word problems (Roy and Roth, 2015). The statistics of these datasets are summarized in Table 1. This comprehensive evaluation approach ensures that the SLMs’ mathematical reasoning capabilities are thoroughly tested across a variety of problem types and structures, providing a robust assessment of their performance.
4.2 Implementation
In our paper, we use GPT-4 as our teacher LLM to generate reasoning processes for questions. Specifically, we manually construct a demonstration set containing eight demonstrations. We then use this demonstration set to prompt GPT-4 to generate four reasoning paths for each question. Each reasoning path is subsequently filtered, resulting in the creation of a KPDD dataset. Next, we use FlanT5 models—Small (60M), Base (250M), and Large (760M) (Chung et al., 2022)—as our student LMs. By fine-tuning the FlanT5 models with the KPDD dataset, we aim to enhance their mathematical reasoning abilities. During the fine-tuning process, we set the learning rate to 5e-4, the batch size to 32, and the total number of training epochs to 10.
Models #Params GSM8K ASDiv SVAMP MultiArith AVG
Proprietary Large Language Models
GPT-4 (OpenAI, 2023) - 92.0 91.3 93.1 - 92.13
ChatGPT - 80.8 87.3 83.0 - 83.7
Claude-2 (Anthropic, 2023) - 85.2 - - - 85.2
PaLM-2 (Anil et al., 2023) 540B 80.7 - - - 80.7
Open-Source Large Language Models
Llama-2 (Touvron et al., 2023) 7B 13.3 50.7 38.0 - 34
CodeLLaMA (Rozière et al., 2023) 7B 34.0 61.4 59.0 - 51.46
Platypus-2 (Lee et al., 2023) 7B 14.4 47.9 36.7 - 33
WizardMath (Luo et al., 2023) 7B 54.9 59.1 57.3 - 57.1
TORA (Gou et al., 2023) 7B 68.8 73.9 68.2 - 70.3
Fine-tuned Small Language Models
Ho et al. Ho et al. (2023) 0.3B 3.11 - - - 3.11
Fu et al. Fu et al. (2023) 0.76B 20.2 23.8 20.4 38.5 25.72
Fu et al. Fu et al. (2023) 0.25B 13.4 20.9 14.2 29.7 19.55
Shridhar et al. Shridhar et al. (2023) 0.77B 17.89 - 18.14 - 18.01
Zhu et al. Zhu et al. (2023) 0.77B 39.2 51.2 48.2 79.2 54.45
Zhu et al. Zhu et al. (2024) 0.77B 42.45 52.81 49.59 85.5 57.58
Our fine-tuned Small Language Models
FlanT5-Small 0.06B 2.1 2.8 2.1 4.0 2.75
(+) KPDD-CoT 7.58 8.73 6.9 7.83 7.76
(+) KPDD-PoT 20.77 40.07 34.1 44.16 34.93
FlanT5-Base 0.25B 3.0 4.2 3.8 7.0 4.5
(+) KPDD-CoT 14.63 14.93 13.8 21.5 16.21
(+) KPDD-PoT 34.57 52.29 50.5 73.66 52.75
FlanT5-Large 0.76B 6.9 10.1 6.8 13.0 9.2
(+) KPDD-CoT 21.75 22.51 19.1 35.5 24.71
(+) KPDD-PoT 46.32 59.92 61.6 87.5 63.83
Table 2: Overall test set performance.
4.3 Main Results
Table 2 showcases our method’s performance on four mathematical datasets, revealing key insights:
  1. 1.
    KPDD-CoT Enhances Mathematical Reasoning: When FlanT5-small is used as the SLM, KPDD-CoT achieves an average accuracy improvement of 5.01% across several mathematical reasoning tasks. With FlanT5-base as the SLM, KPDD-CoT yields an average accuracy improvement of 11.71%. For FlanT5-large, KPDD-CoT results in an average accuracy improvement of 15.51%. The experimental result demonstrates that KPDD-CoT can significantly enhance the mathematical reasoning performance of SLMs. We attribute this experimental result to the reason that baselines often encounters semantic misunderstanding errors that hinder the improvement of SLMs’ mathematical reasoning abilities. In contrast, KPDD-CoT employs extra SLMs to extract key points (including the core question and problem-solving information) of the question and uses these key points to guide the SLMs’ reasoning. This approach significantly reduces the semantic misunderstanding errors of CoTD, making KPDD-CoT better suited for improving the mathematical reasoning ability of SLMs.
  2. 2.
    KPDD-PoT Outperforms State-of-the-Art: When FlanT5-small is used as the SLM, KPDD-PoT achieves an average accuracy improvement of 32.18% across several mathematical reasoning tasks. With FlanT5-base as the SLM, KPDD-PoT yields an average accuracy improvement of 48.25%. For FlanT5-large, KPDD-PoT results in an average accuracy improvement of 54.63%. The experimental result shows that KPDD-PoT make SLMs achieve state-of-the-art mathematical reasoning accuracy. Furthermore, KPDD-PoT’s accuracy is higher than that of KPDD-CoT, highlighting the advantage of rationales in PoT format in enhancing SLMs’ reasoning capabilities. Our analysis finds that the mathematical reasoning performance of CoTD is limited not only by semantic misunderstanding errors but also by calculation errors. PoTD converts rationales from CoT format into PoT format, formulating the reasoning process into a Python program and sending it to an extra Python interpreter to generate the final answer. This method transfers numerical computation from SLMs to a Python interpreter, avoiding calculation errors. Additionally, by extracting key points of the question, KPDD-PoT implicitly enhances the SLMs’ understanding of the question, thereby improving their overall mathematical reasoning capabilities.
  3. 3.
    Strong Transferability of KPDD: KPDD exhibits strong transferability. The distillation dataset of KPDD is constructed based on the GSM8K training dataset, and we evaluate our SLMs on several mathematical reasoning datasets, including the GSM8K test dataset, ASDiv dataset, SVAMP dataset, and MultiArith dataset. Our experimental results show that KPDD not only achieves good reasoning performance on the GSM8K test dataset but also performs well on the ASDiv, SVAMP, and MultiArith datasets. These results demonstrate that KPDD has strong transferability and further corroborate that SLMs do not improve their reasoning performance through data leakage.
Category Core Info Solve GSM8K ASDiv SVAMP MultiArith AVG
1 $\times$ $\times$ $\times$ 3.0 4.2 3.8 7.0 4.5
2 $\times$ $\times$ $\checkmark$ 8.71 9.2 8.2 10.33 9.11
3 $\checkmark$ $\times$ $\checkmark$ 9.02 9.25 8.9 11.5 9.66
4 $\times$ $\checkmark$ $\checkmark$ 8.87 9.73 8.9 11.0 9.59
5 $\checkmark$ $\checkmark$ $\checkmark$ 9.17 9.92 9.03 11.83 9.98
Table 3: Effect of Different Components in KPDD-CoT. We consider five different categories to analyse the effect of different components in KPDD-CoT. The experiment result shows that key points in questions can deepen SLMs’ understanding of the questions, and combining several key points can provide richer information, leading to further improvements in SLMs’ reasoning abilities.
Category Core Info Solve GSM8K ASDiv SVAMP MultiArith AVG
1 $\times$ $\times$ $\times$ 3.0 4.2 3.8 7.0 4.5
2 $\times$ $\times$ $\checkmark$ 19.40 44.32 40.6 45.33 37.41
3 $\checkmark$ $\times$ $\checkmark$ 23.19 45.89 44.1 53.33 41.62
4 $\times$ $\checkmark$ $\checkmark$ 25.39 46.85 44.6 57.33 43.54
5 $\checkmark$ $\checkmark$ $\checkmark$ 27.06 49.33 46.1 58.33 45.20
Table 4: Effect of Different Components in KPDD-PoT. We consider five different categories to analyse the effect of different components in KPDD-PoT. The experiment result shows that key points in questions can deepen SLMs’ understanding of the questions, and combining several key points can provide richer information, leading to further improvements in SLMs’ reasoning abilities.
4.4 Effect of Different Components in KPDD
In this subsection, we delve into the impact of various components within KPDD. We have considered five distinct categories, which include: 1. Original SLMs without any fine-tuning; 2. SLMs with original CoT/PoT distillation; 3. SLMs with core distillation combined with CoT/PoT distillation; 4. SLMs with problem-solving information distillation combined with CoT/PoT distillation; 5. SLMs with KPDD. For each of the latter four categories, we have constructed corresponding reasoning datasets, each containing a single reasoning path per question. Following this, we have utilized FlanT5-base as our foundation for SLMs, and we have fine-tuned these models using the aforementioned reasoning datasets. To evaluate the reasoning capabilities of these SLMs, we have tested them on the GSM8K test dataset, as well as on the ASDiv, SVAMP, and MultiArith datasets.
Tables 3 and 4 present the results of our experiments, from which we make several observations: (1) We observe a significant performance improvement in Category 2 compared to original SLMs. Specifically, under CoT reasoning, Category 2 achieves an average accuracy gain of 4.61% across multiple datasets, while under PoT reasoning, it achieves a substantial average accuracy improvement of 32.91%. These experimental results indicate that CoTD and PoTD can markedly enhance the mathematical reasoning ability of SLMs. (2) We find that Categories 3 and 4 exhibit a further performance increase relative to Category 2. Specifically, in the context of CoT reasoning, Categories 3 and 4 achieve average accuracy gains of 0.55% and 0.45% respectively over Category 2 across multiple datasets. Under PoT reasoning, the gains are more pronounced with Categories 3 and 4 achieving average accuracy improvements of 4.21% and 6.13% respectively. This suggests that SLMs can deepen their understanding of questions by focusing on key points, thereby further enhancing their mathematical reasoning ability. (3) In Category 5, we combine the core questions with the problem-solving information to guide SLMs in addressing the questions. The results are promising: Category 5 achieves an average accuracy of 9.98% under CoT reasoning and a remarkable 45.20% under PoT reasoning across multiple datasets. This indicates that key points in questions play a crucial role in boosting the reasoning capabilities of SLMs, and that combining several key points provides richer information, leading to further improvements in their reasoning abilities.
Category Core Info Solve GSM8K ASDiv SVAMP MultiArith AVG
$\mathrm{I}$ 1* 1 1 7.88 4.72 5.4 10.66 7.16
$\mathrm{II}$ 1 1 2 9.09 9.44 8.2 11.33 9.51
$\mathrm{III}$ 1 2 2 8.41 7.72 6.7 11.24 8.51
$\mathrm{IV}$ 2 1 2 7.80 7.58 7.1 11.16 8.41
$\mathrm{V}$ 1 2 3 9.17 9.92 9.03 11.83 9.98
    The index of SLM.
Table 5: Effect of SLM Quantity in KPDD-CoT. We consider five different categories to analyse the effect of SLM quantity in KPDD-CoT. The experimental results show that for KPDD-CoT, using a separate SLM for each component is necessary to maximize the reasoning performance of KPDD-CoT.
Category Core Info Solve GSM8K ASDiv SVAMP MultiArith AVG
$\mathrm{I}$ 1* 1 1 24.18 44.32 41.19 48.66 39.61
$\mathrm{II}$ 1 1 2 26.0 42.69 42.69 55.83 41.80
$\mathrm{III}$ 1 2 2 24.79 46.37 40.6 49.16 40.23
$\mathrm{IV}$ 2 1 2 24.63 45.37 41.3 49.33 40.15
$\mathrm{V}$ 1 2 3 27.06 49.33 46.1 58.33 45.20
    The index of SLM.
Table 6: Effect of SLM Quantity in KPDD-PoT. We consider five different categories to analyse the effect of SLM quantity in KPDD-PoT. The experimental results show that for KPDD-PoT, using a separate SLM for each component is necessary to maximize the reasoning performance of KPDD-PoT.
4.5 Effect of SLM Quantity in KPDD
In this subsection, we investigate the impact of SLM Quantity in KPDD. We consider five distinct categories: $\mathrm{I}$. Using one SLM to simultaneously extract the core question and problem-solving information, and solve the original question; $\mathrm{II}$. Using one SLM to extract the core question and problem-solving information, and another SLM to solve the original question; $\mathrm{III}$. Using one SLM to extract the core question, another SLM to extract the problem-solving information, and a third SLM to solve the original question; $\mathrm{IV}$. Using one SLM to extract the problem-solving information, another SLM to extract the core question, and both to solve the original question; $\mathrm{V}$. Using one SLM to extract the core question, another SLM to extract the problem-solving information, and a third SLM to solve the original question. For each category, we create corresponding reasoning datasets, each containing a single reasoning path per question. We utilize FlanT5-base as our base SLMs, fine-tuning them on these reasoning datasets. To assess their reasoning capabilities, we evaluate these SLMs on the GSM8K test dataset, as well as on the ASDiv, SVAMP, and MultiArith datasets.
Table 5 and 6 present the results of our experiments, from which we make several observations: (1) Compared to other categories, Category $\mathrm{I}$ performed worse. For KPDD-CoT, Category $\mathrm{I}$ achieved an average accuracy of 7.16% across multiple datasets, while for KPDD-PoT, it achieved an average accuracy of 39.61%. This suggests that the limited model size of a single SLM hinders its performance across multiple tasks. (2) Category $\mathrm{II}$ outperformed Categories $\mathrm{III}$ and $\mathrm{IV}$ in reasoning performance. For KPDD-CoT, Category $\mathrm{II}$ achieved an average accuracy of 9.51% across multiple datasets, while for KPDD-PoT, it achieved an average accuracy of 41.80%. We attribute this result to the importance of the KPDD-CoT/PoT-solve component, where using a single SLM for this phase yields the best reasoning performance. (3) For KPDD-CoT, Category $\mathrm{V}$ achieved an average accuracy of 9.98% across multiple datasets, while for KPDD-PoT, it achieved an average accuracy of 45.20%. This is the highest reasoning performance among all categories, indicating that our approach of using a separate SLM for each component maximizes the performance of each component, thereby maximizing the reasoning performance of KPDD.
Figure 5: Effect of Reasoning Paths. We fine-tune CodeT5-Base with different reasoning paths to analyse the effect of reasoning paths. The experiment results shows that diverse reasoning paths can improve SLMs’ reasoning performance.
4.6 Diverse Reasoning Paths Improve SLMs’ Reasoning Performance
In this subsection, we primarily explore the impact of the diversity of reasoning paths on the mathematical reasoning performance of SLMs. Specifically, we fine-tune FlanT5-base using KPDD datasets with varying numbers of reasoning paths and then evaluate the fine-tuned SLMs on several mathematical reasoning datasets. By analyzing their mathematical reasoning performance, we assess how the diversity of reasoning paths affects the SLMs’ capabilities in mathematical reasoning.
Figure 6: Error Analysis for SLMs. We conducted an error analysis of four different categories of distillation methods. The experiment results show that integrating multiple key points of the questions can significantly reduce SLMs’ understanding errors, enhance the comprehension of the questions and further improve the reasoning performance of SLMs.
Figure 5 reveals the relevant experimental results. As the number of reasoning paths in the distillation dataset increases, the SLM tends to achieve better performance. For example, when fine-tuning FlanT5-base with the KPDD-CoT dataset containing a single reasoning path, it achieved an accuracy of 9.17% on the GSM8K test dataset, 9.92% on ASDiv, 9.03% on SVAMP, and 11.83% on MultiArith. However, when fine-tuning FlanT5-base with the KPDD-CoT dataset containing four reasoning paths, it achieved accuracies of 14.63% on the GSM8K test dataset, 14.93% on ASDiv, 13.8% on SVAMP, and 21.5% on MultiArith. Similarly, when fine-tuning SLMs with KPDD-PoT datasets containing different reasoning paths, the experimental results were consistent with those of KPDD-CoT. These results indicate that a mathematical reasoning dataset with diverse reasoning paths can effectively enhance the mathematical reasoning performance of SLMs.
4.7 Error Analysis
In this subsection, our aim is to verify whether KPDD can indeed reduce semantic misunderstanding errors. KPDD-PoT implicitly includes the reasoning process within its rationales, making it challenging to conduct error analysis on rationales in PoT format. Conversely, rationales in CoT format explicitly contain the reasoning steps, allowing us to clearly understand how the SLM solves the questions step by step, thus facilitating error analysis. Therefore, in this part, we focus on error analysis for rationales in CoT format. To achieve our goal, we randomly sample 100 examples from GSM8K/SVAMP and perform error analysis on the questions with incorrect answers. For a better understanding of KPDD’s effect, we also consider three other scenarios: (1) vanilla CoTD, (2) reasoning that combines vanilla CoTD and core question extraction, and (3) reasoning that combines vanilla CoTD and problem-solving information extraction. Furthermore, to simplify our analysis, we use flanT5-base as our SLMs, and the corresponding reasoning datasets still contain a single reasoning path per question.
The detailed quantitative results are illustrated in Figure 6. By analyzing the experimental results, we found that: (1) Combination of Multiple Errors in SLMs: SLMs tend to exhibit combinations of multiple errors, with calculation errors having the most significant impact on reasoning performance. Specifically, vanilla CoTD on the GSM8K dataset showed 51 understanding errors, 79 calculation errors, and 34 step missing errors, resulting in a total of 164 errors. This number far exceeds the original number of problems, with calculation errors outnumbering other types of errors. Similar results were observed in the SVAMP dataset. This explains why PoTD achieves better reasoning performance than CoTD: PoTD converts vanilla rationales into Python programs, delegating the calculation process to an external Python interpreter to avoid calculation errors. (2) Reduction of Understanding Errors with Key Points: Introducing key points of the original questions effectively reduces understanding errors. Specifically, when core questions were introduced in vanilla CoTD, the number of understanding errors on the GSM8K dataset decreased to 50, and on the SVAMP dataset, it decreased to 53. When problem-solving information was introduced in vanilla CoTD, the number of understanding errors decreased to 48 on GSM8K and to 51 on SVAMP. These results indicate that key points of the original questions help SLMs better understand the questions, thereby reducing understanding errors and improving reasoning performance. (3) Further Reduction of Understanding Errors with Multiple Key Points: Combining multiple key points can further reduce understanding errors. Specifically, KPDD reduced the number of understanding errors to 46 on GSM8K and to 50 on SVAMP. This suggests that KPDD’s method of integrating multiple key points can deepen SLMs’ understanding of the original questions, further reducing understanding errors and enhancing reasoning performance.
5 Conclusion
In this paper, we propose Key-Point-Driven Distillation (KPDD) for enhancing mathematical reasoning in Small Language Models (SLMs). Our approach leverages the extraction of key points from questions to improve understanding and reduce errors in reasoning tasks. Experimental results demonstrate that KPDD significantly reduces understanding errors compared to conventional mathematical reasoning distillation method. However, PoTD implicitly embeds the reasoning process within the generated program, making it difficult to analyze misunderstandings. In the future, we will explore error analysis methods to facilitate PoTD error analysis.